OpenStudy (anonymous):

the sum of 2 #'s is 141. one # is 37 more than the other. FIND THE NUMBERS

OpenStudy (anonymous):

89 and 52

OpenStudy (anonymous):

THANKU

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

let one number=x other number=x+37 so,x+37+x=141 2x=104 x=52 other number=37+52=89